# Freeze-Omni: A Smart and Low Latency Speech-to-speech Dialogue Model with Frozen LLM

### v2

@@ Abstract
Rapidly developing large language models (LLMs) have brought tremendous intelligent applications. GPT-4o’s excellent duplex speech interaction ability has recently brought impressive experience to users. Researchers have recently proposed several multi-modal LLMs in this direction that can achieve speech-to-speech dialogue. This paper proposes a novel speech-text multimodal LLM architecture called Freeze-Omni. Our main contribution is that the speech input and output modalities can be easily connected to a textual LLM while keeping the LLM’s parameters frozen throughout the training process. We designed 3-stage training strategies both for the modeling of speech input and output, enabling Freeze-Omni to obtain speech-to-speech dialogue ability using text-speech paired data (such as ASR and TTS data) and only 60,000 multi-round text Q&A data on 8 GPUs. Moreover, we can effectively ensure that the intelligence of the Freeze-Omni in the speech modality is at the same level compared with that in the text modality of its backbone LLM, while the end-to-end latency of the spoken response achieves a low level. In addition, we also designed a method to achieve duplex dialogue ability through multi-task training, making Freeze-Omni have a more natural style of dialogue ability between the users. Freeze-Omni mainly provides a possibility for researchers to conduct multimodal LLM under the condition of a frozen LLM, avoiding various impacts caused by the catastrophic forgetting of LLM caused by fewer data and training resources.
@@ 1 Introduction
In recent years, the development of large language models has been extremely rapid. A series of large language models represented by the GPT series [10, 1] of OpenAI has demonstrated extraordinary capabilities. As speech interaction is one of the most natural forms of human-computer interaction, combining speech input and output with an LLM can bring an extraordinary experience to users. The traditional method is to use a cascaded approach of ASR + LLM + TTS to achieve the interaction with LLM in speech modality. However, this approach often leads to a relatively high engineering complexity and a considerable interaction latency. Nevertheless, GPT-4o [18] has changed this situation, it provides an end-to-end speech interaction mode which has significantly improved the user experience, triggering a research boom among researchers regarding multimodal LLMs for speech-to-speech interaction.
In the field of general LLMs, many public models such as Llama 3.2 [8], Qwen2.5 [21], Mixtral [14], etc. have provided very good opportunities for researchers to develop downstream tasks on them. Therefore, in the research field of multimodal LLMs for speech-to-speech, works such as Mini-Omni2 [24], LLaMA-Omni [9], and Moshi [7] have provided excellent references for researchers. These works adopt different strategies to align the speech modality with the LLM and design some methods to achieve a duplex dialogue mode, demonstrating excellent performance.
In this research context, we found that in the process of aligning the LLM with the speech modality in existing public speech-text multimodal LLMs [6, 7, 9, 11, 27, 23], the parameters of the LLM are more or less fine-tuned. However, in most cases, it is very difficult for researchers to easily collect spoken Q&A data at the million-hour level (the corresponding text content can be comparable to the amount of data for training text-modal LLMs). This inevitably brings about the forgetting problem to the LLM, resulting in a negative impact on its intelligence. In addition, only a few works have evaluated the accuracy of spoken question-answering tasks for speech-to-speech multimodal LLMs, and show an obvious gap in performance between spoken question-answering and text-modality question-answering. Therefore, in this paper, we propose a speech-to-speech dialogue LLM called Freeze-Omni, achieving speech modality alignment while the LLM is frozen throughout the training process, and obtaining low latency speech dialogue capabilities while keeping the intelligence of the backbone LLM. Freeze-Omni is mainly implemented in the following steps:
@@
Modeling of speech output Second, we use a mount of text-speech paired data to train the AR-based speech decoder which can generate speech tokens from text and a single-codebook based codec model is used to decode the speech token into waveform. Then, we design a prefix kv-cache fine-tune strategy, using the hidden state vector output by the LLM to transfer the speech decoder into the output text space of LLM, achieving the ability of text input to speech output while keeping the LLM frozen.
Design for duplex dialogue Finally, we simultaneously connect the speech encoder and speech decoder from the above parts to the backbone LLM. Then, a task of chunk-wise state prediction is used to determine whether or not the user interrupts the dialogue, achieving the duplex speech-to-speech dialogue ability.
In conclusion, the main contributions of the proposed Freeze-Omni are as follows:
  • The parameters of the LLM are completely frozen throughout the training process, ensuring that the intelligence of the LLM will be kept. At the same time, the ability of low latency speech-to-speech dialogue is still obtained.
  • The data scale relied on during the training process is small and consumes fewer computing resources. It requires text-speech paired data (such as ASR and TTS training data) and only a small amount of Q&A data in text modality.
@@ 3.1.3 Training
3.2 Results on speech input
3.3 Results on speech output
@@ 4 Conclusion and future work
In this paper, we proposed Freeze-Omni, a text-audio multimodal LLM capable of low-latency speech-to-speech dialogue, which does not need fine-tuning the LLM, showing excellent performance in various evaluation tasks. In the future, to explore more speech dialogue capabilities, we plan to do the following work:
  • We will upgrade the speech encoder to an audio encoder so that it can process and understand non-speech to complete tasks such as emotion understanding and audio captioning.
  • We will add more multi-tasks to make the LLM output more task labels to complete more downstream tasks of speech dialogue, taking the state prediction ability as an example, under the condition of LLM freeze.
  • We will explore how to support multi-speaker synthesis and instruct follow ability in the speech decoder part so that it can obtain more instruct information from the hidden state of the LLM and provide more abundant speech output styles.

### v3

@@ Abstract
Rapidly developing large language models (LLMs) have brought tremendous intelligent applications. Especially, the GPT-4o’s excellent duplex speech interaction ability has brought impressive experience to users. Researchers have recently proposed several multi-modal LLMs in this direction that can achieve user-agent speech-to-speech conversations. This paper proposes a novel speech-text multimodal LLM architecture called Freeze-Omni. Our main contribution is that the speech input and output modalities can be easily connected to a textual LLM while keeping the LLM’s parameters frozen throughout the training process. We design a three-stage training strategy for modeling both the speech input and output, enabling Freeze-Omni to obtain speech-to-speech conversation ability using text-speech paired data (such as ASR and TTS data) and only 60,000 multi-round text Q&A data on 8 GPUs. Moreover, we can effectively ensure that the intelligence of the Freeze-Omni in the speech modality is at the same level compared with that in the text modality of its backbone LLM, while achieving low latency end-to-end spoken response. In addition, we also designed a method to achieve duplex dialogue ability through multi-task training, giving Freeze-Omni a more natural style of dialogue ability between users and agents. In summary, Freeze-Omni holds great potential to conduct speech-to-speech dialogue based on a multimodal LLM under the condition of a frozen LLM, avoiding the catastrophic forgetting problem caused by limited data and training resources.
@@ 1 Introduction
Recent years have witnessed rapid development of large language models (LLMs). The family of LLMs represented by the GPT series [10, 1] of OpenAI has demonstrated extraordinary capabilities. As speech interaction is one of the most natural forms of human-computer interaction, combining speech input and output with an LLM can bring a natural experience to users. The traditional method is to use a cascaded approach of ASR + LLM + TTS to achieve the interaction with LLM in speech modality. However, this engineering-centered pipeline approach often leads to a considerable interaction latency. Nevertheless, GPT-4o [18] has changed this situation – it provides an end-to-end speech interaction mode which has significantly improved user experience, triggering a research boom regarding multimodal LLMs for speech-to-speech interaction.
In the field of general LLMs, many public models such as Llama 3.2 [8], Qwen2.5 [21], Mixtral [14], etc., have provided good opportunities for researchers to develop downstream tasks. Therefore, in the field of multimodal LLMs targeting speech-to-speech conversation, works such as Mini-Omni2 [24], LLaMA-Omni [9], and Moshi [7] have provided excellent references for researchers. These works adopt different strategies to align the speech modality with an LLM and design some tricks to achieve a duplex dialogue mode.
In this research context, we found that in the process of aligning the LLM with the speech modality in existing public speech-text multimodal LLMs [6, 7, 9, 11, 27, 23], the parameters of the LLM are more or less fine-tuned. However, in most cases, it is very difficult for researchers to easily collect spoken Q&A data at the million-hour level (the corresponding text content can be comparable to the amount of data for training text-modal LLMs). Fine-tuing the LLM inevitably brings about the catastrophic forgetting problem to the LLM, resulting in a negative impact on its original “intelligence” ability. In addition, only a few works have evaluated the perfomrance of spoken question-answering tasks for speech-to-speech multimodal LLMs, showing an obvious gap in the performance between spoken Q&A and text-modality Q&A. Therefore, in this paper, we propose a speech-to-speech dialogue LLM called Freeze-Omni, which achieves effective speech-text modality alignment while keeping the LLM parameters frozen, obtaining low latency speech dialogue capabilities while maintaining the orignal intelligence of the backbone LLM. Specifically, Freeze-Omni is mainly implemented in the following steps:
@@
Modeling of speech output Second, we use a sizable amount of text-speech paired data to train the AR-based speech decoder which can generate speech tokens from text and a single-codebook based codec model is used to decode the speech token into waveform. Then, we design a prefix kv-cache fine-tune strategy, using the hidden state vector output by the LLM to transfer the speech decoder into the output text space of LLM, achieving the ability of text input to speech output while keeping the LLM frozen.
Design for duplex dialogue Finally, we connect the speech encoder and speech decoder from the above parts to the backbone LLM. Then, a task of chunk-wise state prediction is used to determine whether or not the user interrupts the dialogue, achieving the duplex speech-to-speech dialogue ability.
In conclusion, the main contributions of Freeze-Omni are as follows:
  • The parameters of the LLM are completely frozen throughout the training process of Freeze-Omni, maintaining the original intelligence of the LLM and achieving low latency speech-to-speech dialogue at the same time.
  • The paired text-speech Q&A training data is at a small scale and consumes fewer computing resources in the building of Freeze-Omni.
@@ 3.1.3 Training
3.2 Results on speech input
3.3 Results on speech output
In addition, the NAR and AR speech decoders need to model the LLM embedding outputs and speech tokens simultaneously, but the spaces represented by these two are different. Therefore, to verify whether the generation quality would be improved if the NAR speech decoder has additional parameters for modeling the outputs of the LLM embedding layer compared to the AR speech decoder, we added an extra pre-network between the NAR speech decoder and the LLM embedding layer. This pre-network consists of two Llama decoder layers with the same configuration of the NAR speech decoder. As shown in Tab 2, this method can significantly improve the speech quality generated by the speech decoder.
@@ 4 Conclusion and future work
In this paper, we proposed Freeze-Omni, a text-audio multimodal LLM capable of low-latency speech-to-speech dialogue, which does not need fine-tuning the LLM backbone, showing excellent performance in various tasks. In the future, to explore more speech dialogue capabilities, we plan to do the following updates:
  • We will upgrade the speech encoder to a general audio encoder so that it can complete tasks such as emotion understanding and audio captioning.
  • Under the condition of a frozen LLM, we will add more tasks to make the LLM complete more downstream tasks in speech dialogue like the state prediction ability.
  • We plan to support multiple voices and instruct-following ability in the speech decoder part so that it can obtain more instruct information from the hidden state of the LLM and provide richer speaking styles.